# Web-CogReasoner: Towards Knowledge-Induced Cognitive Reasoning for Web Agents

### v1

@@ A.2.2 Understanding Web Knowledge
  • Embedded Image QA: Given a question grounded in the content of an embedded image within a webpage screenshot, the model must produce an accurate, context-aware answer using only visual information. These questions may refer to image content (e.g., "What brand is shown in the ad?") or its function in the UI.
@@
All four subtasks are derived from the Multi-UI [23] dataset, which provides rich annotations for webpage visual elements and user-facing semantics. Together, these subtasks measure a model’s ability to perform grounded visual-language understanding at both local (element-level) and global (page-level) scales.
A.2.3 Procedual Web Knowledge
User’s intention Pre
@@ A.2.3 Procedual Web Knowledge
    Task Definition: Given a sequence of web page screenshots $p_{1},p_{2},\ldots,p_{n}$ representing a user’s interaction trajectory, the objective is to predict the original user instruction y that guided the sequence. Each screenshot $p_{t}$ corresponds to the visual observation at step $t$ of a successful task execution. Formally, the model learns a mapping: $f:\{p_{1},p_{2},……,p_{n}\}\rightarrow Q$ where Q is the natural language instruction.
@@
We curated a collection of 51 popup components from JS Design website, encompassing a diverse range of visual styles and functional categories, such as notification modals, alert dialogs, and login forms. This diversity ensures comprehensive coverage of real-world popup use cases. For background webpages, we utilized the OpenWebVoyager[13] dataset, which contains a large number of authentic webpage screenshots with varied layouts and content, providing a rich foundation for synthesizing realistic popup-injected webpages. To construct the training data for this task, we employed the following procedure:
@@
    Generating Popup AX Tree: Each popup image was processed using Qwen-VL 2.5 32B to generate an ARIA-compliant AX Tree. To simulate diverse structural configurations, we randomly modified the index values of popup AX Tree elements and inserted the popup AX Tree into different locations within the original webpage’s AX Tree, resulting in a combined AX Tree that reflects realistic variations in webpage structure.
@@
    Generating Popup Closing Strategies: We then instructed Qwen-VL 2.5 32B to identify all n possible methods for closing the popup, based on the popup image and its corresponding AX Tree. Recognizing that, in practical settings, any correct method is sufficient, we applied combinatorial augmentation to the n methods. Specifically, we enumerated all non-empty subsets of the n strategies, yielding a total of $2^{n}-1$ distinct answer combinations. This expansion significantly broadens the training distribution and increases the model’s exposure to diverse correct solutions.
@@
We define the Single-Step Web Task to evaluate a model’s ability to ground high-level user intentions in visual webpage elements. Each task instance includes a full-page screenshot from a real-world webpage, a concise natural language instruction (e.g., "Search for a product", "Log into the system"), and several candidate elements marked by red bounding boxes.
@@
All samples are directly sourced from the Multi-UI [23] dataset, which provides rich, annotated webpage screenshots paired with task descriptions and labeled ground-truth targets. No trajectory-level annotation or external instruction rewriting is involved. This task offers a reliable benchmark for evaluating atomic web interaction capabilities in a static, visually grounded setting.
@@
To further enhance the original OpenWebVoyager [13] dataset, we incorporate Knowledge-driven Chain-of-Thought (CoT) Reasoning to improve the model’s stepwise understanding and execution, details see Section  4.2 In addition, to simulate realistic interruptions during multi-step web interactions, we propose the Noisy Multi-Step Web Task by augmenting interaction trajectories from OpenWebVoyager. Specifically, for each sample in our Popup Close dataset, a popup window is injected at a specific step (e.g., step $t$) of an existing task trajectory.

### v2

@@ A.2.4 Understanding Web Knowledge
  • Embedded Image QA: Given a question grounded in the content of an embedded image within a webpage screenshot, the model must produce an accurate, context-aware answer using only visual information. These questions may refer to image content (e.g., ”What brand is shown in the ad?”) or its function in the UI.
@@
All four subtasks are derived from the Multi-UI (Liu et al., 2024a) dataset, which provides rich annotations for webpage visual elements and user-facing semantics. Together, these subtasks measure a model’s ability to perform grounded visual-language understanding at both local (element-level) and global (page-level) scales.
A.2.5 Procedural Web Knowledge
User’s intention Prediction
@@ A.2.5 Procedural Web Knowledge
    Task Definition: Given a sequence of web page screenshots $p_{1},p_{2},\ldots,p_{n}$ representing a user’s interaction trajectory, the objective is to predict the original user instruction y that guided the sequence. Each screenshot $p_{t}$ corresponds to the visual observation at step $t$ of a successful task execution. Formally, the model learns a mapping: $f:\{p_{1},p_{2},\ldots,p_{n}\}\rightarrow Q$ where $Q$ is the natural language instruction.
@@
We curated a collection of 51 popup components from JS Design website, encompassing a diverse range of visual styles and functional categories, such as notification modals, alert dialogs, and login forms. This diversity ensures comprehensive coverage of real-world popup use cases. For background webpages, we utilized the OpenWebVoyager(He et al., 2024b) dataset, which contains a large number of authentic webpage screenshots with varied layouts and content, providing a rich foundation for synthesizing realistic popup-injected webpages. To construct the training data for this task, we employed the following procedure:
@@
    Generating Popup AX Tree: Each popup image was processed using Qwen-VL-2.5-32B to generate an ARIA-compliant AX Tree. To simulate diverse structural configurations, we randomly modified the index values of popup AX Tree elements and inserted the popup AX Tree into different locations within the original webpage’s AX Tree, resulting in a combined AX Tree that reflects realistic variations in webpage structure.
@@
    Generating Popup Closing Strategies: We then instructed Qwen-VL-2.5-32B to identify all n possible methods for closing the popup, based on the popup image and its corresponding AX Tree. Recognizing that, in practical settings, any correct method is sufficient, we applied combinatorial augmentation to the n methods. Specifically, we enumerated all non-empty subsets of the n strategies, yielding a total of $2^{n}-1$ distinct answer combinations. This expansion significantly broadens the training distribution and increases the model’s exposure to diverse correct solutions.
@@
We define the Single-Step Web Task to evaluate a model’s ability to ground high-level user intentions in visual webpage elements. Each task instance includes a full-page screenshot from a real-world webpage, a concise natural language instruction (e.g., ”Search for a product”, ”Log into the system”), and several candidate elements marked by red bounding boxes.
@@
All samples are directly sourced from the Multi-UI (Liu et al., 2024a) dataset, which provides rich, annotated webpage screenshots paired with task descriptions and labeled ground-truth targets. No trajectory-level annotation or external instruction rewriting is involved. This task offers a reliable benchmark for evaluating atomic web interaction capabilities in a static, visually grounded setting.
@@
To further enhance the original OpenWebVoyager (He et al., 2024b) dataset, we incorporate Knowledge-driven Chain-of-Thought (CoT) Reasoning to improve the model’s stepwise understanding and execution, details see Section  4.2 In addition, to simulate realistic interruptions during multi-step web interactions, we propose the Noisy Multi-Step Web Task by augmenting interaction trajectories from OpenWebVoyager. Specifically, for each sample in our Popup Close dataset, a popup window is injected at a specific step (e.g., step $t$) of an existing task trajectory.
@@
A.3 Train Details
A.3.1 Training
We employ a multi-phase Imitation Learning strategy to train our model on Web-CogDataset, utilizing Qwen2.5-VL-7B (Bai et al., 2025) as the base model. Each phase is aligned with a distinct layer of the Web-CogKnowledge Framework: (1) the first knowledge content learning focuses on acquiring Factual Knowledge and Conceptual Knowledge, enabling the model to interpret web content and semantics; (2) the second cognitive process emphasizes Procedural Knowledge, training the model to plan and execute multi-step web interactions. To accommodate the increased complexity of the final phase, which involves multi-image inputs and extended reasoning, we configure training with a maximum sequence length of 8K and a batch size of 1 with gradient accumulation of 16 steps. All experiments are conducted on a cluster of 8 x NVIDIA A800 80GB GPUs.
A.3.2 Performance Under Different Training Strategies
We could feed these SFT datasets into models with different training strategies. In this section, We investigate how they influence the model’s performance.
  1. 1.
    Curriculum Learning Strategy: We fine-tune the model following a curriculum learning paradigm.
  2. 2.
    Mixed Multi-task Learning: We directly mix different tasks and apply SFT.
Through comparative analysis, we find that: Under the same task scenario, models trained via curriculum learning conduct multiple rounds of exploration continuously and do not cease exploration prematurely. In contrast, when retrieving task - related data, models trained via mixed training terminate the search within a limited number of attempts if they fail to find results.
As shown in Figure 7, when faced with the user’s instruction ”Open the most helpful 5 - star reviews of Alpine Ridge”, models trained via mixed training deviate from the domain of task - specific information retrieval and autonomously generate an instruction to ”switch to forums for information”.
Based on these observations, we conjecture: Curriculum learning trains models incrementally from simple to complex task, which enables them to develop a fundamental understanding of the problem structure. They continue to carry out multiple rounds of exploration to ensure comprehensive comprehension and accurate processing, thus avoiding the premature termination of exploration. In contrast, mixed training models may be trained on multiple tasks simultaneously. The potential interference among different tasks can make models easily influenced by irrelevant tasks when processing a specific one, thereby undermining their ability to focus on task - specific information retrieval and processing.
Figure 7: Comparison between mixed and curriculum strategies.
A.4 Evaluation Details
A.4.1 LVM Evaluation Protocol
For open-ended generation tasks in Web-CogBench (i.e., Element Understanding, WebPage Understanding, and User Intent Prediction), we utilize a high-capability LVM (GPT-4o) as an automated evaluator. The evaluation process involves a strict comparison between the Candidate Model’s Answer and the verified Ground Truth based on image provided in the dataset.
To ensure fine-grained assessment, we decompose the evaluation into specific cognitive dimensions rather than relying on a single holistic score:
  • Element Understanding: Assessed on Appearance (visual fidelity), Position (structural context), and Function (interaction purpose).
  • WebPage Understanding: Assessed on Structure & Layout, Key Element Analysis, and Summary Coherence.
  • User Intent Prediction: Assessed on Evidence Alignment (visual cue detection), Intent Accuracy, and Reasoning Quality.
We employ a rigorous 1-5 Likert Scale and normalized to 0–100 for scoring. A score of 5 denotes that the candidate ”fully and accurately captures all relevant information present in the Ground Truth,” while lower scores reflect varying degrees of omissions or inaccuracies. The evaluator outputs a structured JSON object containing both integer scores and text justifications for each dimension, ensuring the traceability of the results.
For instance, the Element Understanding task is assessed on:
System Instruction: You are a meticulous and impartial AI evaluator for a web UI understanding benchmark. Your task is to assess the quality of a candidate model’s answer by comparing it strictly against a ground truth and provided image reference. Your evaluation must be based exclusively on the information provided in the “Ground Truth Answer” and ”Image”. Evaluate the candidate answer on three specific aspects: Appearance, Position, and Function. [Ground Truth Answer] {ground_truth} [Candidate Model’s Answer] {model_answer} Evaluation Criteria & Scoring: Score 1: Completely incorrect or missing. Score 2: Mostly incorrect, with a minor element of truth. Score 3: Partially correct, but misses significant details mentioned in the ground truth. Score 4: Mostly correct, with only minor inaccuracies or omissions compared to the ground truth. Score 5: Fully and accurately captures all relevant information present in the ground truth. Your response MUST be a single, valid JSON object, adhering to the following structure. Do not add any text before or after the JSON object. { ”appearance_score”: ¡integer_score¿, ”appearance_justification”: ”¡Your brief justification… referencing the ground truth¿”, ”position_score”: ¡integer_score¿, ”position_justification”: ”¡Your brief justification… referencing the ground truth¿”, ”function_score”: ¡integer_score¿, ”function_justification”: ”¡Your brief justification… referencing the ground truth¿”, ”overall_score”: ¡A final holistic integer score from 1 to 5¿, ”overall_justification”: ”¡A final summary of the model’s performance¿” }
A.4.2 LVM Judge Reliability
To mitigate potential biases from a single evaluator, we employed multiple distinct LVMs (including GPT-4o, Claude Sonnet 4, and Gemini 2.5 Pro) to conduct a rigorous inter-rater reliability analysis. We calculate the ”Within-1-Point Agreement”, defined as the percentage of instances where scores assigned by different LVM judges differ by no more than 1 point.
As shown in Table 14, the high agreement rates across different models confirm that our evaluation criteria are robust and model-agnostic. Furthermore, the strong correlation with Human Proxy Analysis suggests that our Ground-Truth Anchored protocol effectively aligns automated judgment with human evaluation standards.
Table 14: Inter-Rater Reliability Analysis of LVM Judge.
Task Within-1-Point Agreement Human Proxy Analysis
Element Understanding 98.7% 96.7%
WebPage Understanding 97.0% 95.4%
User Intent Prediction 96.0% 94.4%
A.5 Qualitative Analysis
To provide deeper insights into Web-CogReasoner’s capabilities, we present a two-part qualitative analysis. First, we examine the evolution of cognitive abilities across training stages to validate our curriculum. Second, we present a comparative case study on a complex real-world task to demonstrate how our model overcomes knowledge blind spots that trap baseline models.
A.5.1 Evolution of cognitive abilities
Beyond the quantitative improvements shown in our ablation study, a qualitative analysis of the agent’s behavior at each stage offers deeper insights into how our curriculum shapes its cognitive abilities. We examine the agents’ performance on a representative task: ”Find and add a laptop under \$1000 to the shopping cart on an e-commerce website.”
Base Model (Qwen2.5-VL-7B)
Without any specialized training, the base model struggles to formulate a coherent plan. Its reasoning is often generic and untethered from the specific UI. It might correctly identify a ”search bar” but fails to execute a meaningful action, or hallucinates actions that are not possible. For instance, its thought process might be: ”I should search for a laptop,” but its action is an ungrounded ‘click ”Categories” because it lacks the procedural knowledge to connect intent to a multi-step sequence of actions.
Stage 1 Agent (+ Factual Knowledge)
After training on Factual Knowledge, the agent’s perceptual abilities are significantly enhanced. It can now accurately identify and label key elements with their correct attributes. Its thought process becomes grounded in the facts of the page: ”I see a search bar [ID: 25] with the name ’Search products’. I see a button [ID: 28] with the name ’Search’.” However, it still struggles with planning. It understands ”what” is on the page but not ”why” or ”how” to use it. It might correctly type ”laptop” into the search bar but then get stuck, not understanding that the next logical step is to click the search button to submit the query.
Stage 2 Agent (+ Conceptual Knowledge)
With the addition of Conceptual Knowledge, the agent begins to understand the relationships between elements and their purpose. Its reasoning graduates from simple identification to semantic interpretation. The thought process now reflects this understanding: ”The search bar [ID: 25] is for inputting queries. The search button [ID: 28] is functionally linked to it and will trigger the search. This group of elements forms a ’search component’.” This allows it to reliably complete the search and navigate to the result page. However, on the result page, it may still struggle with complex procedural logic, such as applying a price filter.
Full Model (Web-CogReasoner + Procedural Knowledge)
The final agent, equipped with Procedural Knowledge, demonstrates goal-oriented planning and execution. It seamlessly translates the high-level task into a concrete action sequence. Its thought process is now a strategic plan: ”Goal: Add laptop under \$1000 to cart. Step 1: Type ’laptop’ into search bar [ID: 25]. Step 2: Click search button [ID: 28]. Step 3: On the results page, locate the ’Price Range’ filter. Step 4: Input ’1000’ into the ’max price’ field [ID: 57]. Step 5: Identify a suitable product from the filtered list and click its ’Add to Cart’ button [ID: 83].” This demonstrates a complete cognitive loop from perception and understanding to successful action, validating the necessity of the final procedural training stage.
A.5.2 Comparative case study
To validate the necessity of foundational knowledge in handling complex real-world scenarios, we compare the Base Model with Web-CogReasoner on a specific Amazon task.
Task: ”Find a gaming desktop with Windows 11 Home and 1TB disk.”
Base Model Failure: The Knowledge Blind Spot
Lacking explicit knowledge of page layout and element functions, the Base Model literally ”sees” the pixels but ”misses” the affordance.
  • Observation: The model sees the search results but fails to recognize the sidebar filters as the mechanism to refine the query.
  • Error: It misinterprets the page state, assuming a re-search is necessary. It enters a logical dead loop of repeatedly clicking the search button.
  • Action: click [1470] (Search Button) $\to$ Stuck.
Web-CogReasoner Success: Knowledge-Driven Grounding
Leveraging learned Factual and Conceptual knowledge, our agent explicitly identifies the functional role of UI components.
  • Reasoning: The agent identifies the sidebar as a filter section. It conceptually maps the user’s ”1TB” requirement to the specific filter element, predicting that clicking it will narrow the results without leaving the page.
  • Action: click [95] (Filter Link "1 TB") $\to$ Success.
This comparison highlights that success in complex tasks is not just about planning (Procedural), but requires strictly accurate recognition of element functions (Factual/Conceptual) to ground those plans.